OpenStudy (anonymous):

Lisa's living room is 17 feet wide. If the length is 4 feet more than twice the width, what is the area of her living room?

OpenStudy (jim766):

how do you find area of a rectangle?

OpenStudy (jim766):

Area formula for rectangle?

OpenStudy (jim766):

A = L * W look familiar?

OpenStudy (anonymous):

17*38=646

OpenStudy (strawberry17):

w = 17 l = 2w + 4 A = l • w A = (2w + 4) • (w) A = [2(17) + 4] • (17) A = (38) • 17 A = ?